# Lagrangian Skeleta and Koszul Duality on Bionic Symplectic Varieties

### v1

@@ 1. Introduction
Many symplectic varieties that arise in nature can be equipped not only with an elliptic ${\mathbb{G}}_{m}$-action but also a compatible Hamiltonian ${\mathbb{G}}_{m}$-action, i.e. one preserving the symplectic form. When the two actions have the same, finite, set of fixed points, we call the variety $\mathfrak{X}$ bionic symplectic: a precise definition appears in Section 2.2. Standard examples are, again, cotangent bundles of projective spherical varieties, and Hilbert schemes of points on minimal resolutions of cyclic Kleinian singularities (among other quiver varieties). Such a variety comes equipped with a canonical Lagrangian “skeleton” $\mathsf{\Lambda}$.
Suppose now that a symplectic variety $\mathfrak{X}$ comes equipped with a deformation quantization ${\mathcal{W}}$ compatible with an elliptic ${\mathbb{G}}_{m}$-action. Taking ${\mathcal{W}}$-modules equivariant for the elliptic ${\mathbb{G}}_{m}$-action, one arrives at a category useful for geometric representation theory. When the action is the scaling action on fibers of $\mathfrak{X}=T^{*}X$, this category reproduces a category of twisted ${\mathcal{D}}$-modules on $X$; more generally such categories are closely related (depending on the geometry of $\mathfrak{X}$) to representations of Cherednik algebras, finite $W$-algebras, and more. When the variety is bionic symplectic and the deformation quantization is compatible with both ${\mathbb{G}}_{m}$-actions, there is a good geometric analogue of category $\mathcal{O}$ [BLPW1, BLPW2, Lo, Web] inside the category of ${\mathcal{W}}$-modules, consisting of modules supported on the skeleton $\mathsf{\Lambda}$.
The present paper establishes two basic features of the category of ${\mathcal{W}}$-modules on a bionic symplectic variety. We build a sheaf of proper dg algebras $\Omega$ supported (cohomologically) on a Lagrangian $\Lambda\supset\mathsf{\Lambda}$ in $\mathfrak{X}$ that plays the role for ${\mathcal{W}}$-modules analogous to that played by the de Rham complex for ${\mathcal{D}}_{X}$-modules, i.e. ${\mathcal{W}}$-modules on $\mathfrak{X}=T^{*}X$.
@@
Assume that locally free resolutions exist for coherent ${\mathcal{W}}$-modules. Then bounded derived category of coherent ${\mathcal{W}}$-modules is equivalent to the derived category of coherent dg $\Omega$-modules:
@@
As in [BDMN], we define the ${\mathcal{W}}\text{-}{\mathsf{qcoh}}$ to be the ind-category of good ${\mathcal{W}}$-modules. We show that Theorem 1.1 extends to the derived category $D({\mathcal{W}}\text{-}{\mathsf{qcoh}})$, though not in the obvious way. We introduce the ”exotic” derived category $D_{ex}(\Omega\text{-}{\mathsf{Mod}})$ as a certain quotient of $K(\Omega\text{-}{\mathsf{Mod}})$, admitting a quotient functor $D_{ex}(\Omega\text{-}{\mathsf{Mod}})\to D(\Omega\text{-}{\mathsf{Mod}})$ to the usual derived category of dg $\Omega$-modules.
@@ 1.1. Symplectic resolutions
Assume that $\mathfrak{X}\to X$ is a symplectic resolution. Then the sheaf of dg algebras $\Omega$ is cohomologically supported on the Lagrangian skeleton $\mathsf{\Lambda}$.
The key to Theorem 1.4 is the fact that one can choose a local generator $\mathcal{L}$ (whose dg endomorphism ring is $\Omega$) supported on $\mathsf{\Lambda}$; see Proposition 3.16.
Extension of holonomic complexes
@@
${D^{b}_{\mathrm{Hol},C}({\mathcal{W}}_{\mathfrak{X}})}$${D^{b}_{\mathrm{Hol}}({\mathcal{W}}_{\mathfrak{X}})}$${D^{b}_{\mathrm{Hol}}({\mathcal{W}}_{\mathfrak{X}}).}$$\scriptstyle{i_{*}}$$\scriptstyle{\mathbb{R}i^{!}}$$\scriptstyle{i^{*}}$$\scriptstyle{j^{*}}$$\scriptstyle{\mathbb{R}j_{*}}$$\scriptstyle{j_{!}}$
If $S$ denoted the coisotropic reduction of $C$ then it is shown in [BDMN, Theorem 1.7(2)] that the functor of Hamiltonian reduction $\mathbb{H}$ identifies $D^{b}_{\mathrm{Hol},C}({\mathcal{W}}_{\mathfrak{X}})$ with $D^{b}_{\mathrm{Hol}}({\mathcal{W}}_{S})$.
@@ 1.2. Categorical conventions
Funding
@@
Acknowledgements
The authors are grateful to D. Ben-Zvi, K. Kremnitzer, J. Lebowski, D. Nadler, and D. Treumann for inspiration and helpful conversations.
@@ 2.1. Lagrangian Skeletons
    For each $i\in I$, $\Lambda_{i}=\Lambda\cap C_{i}$ is a smooth closed subvariety of $C_{i}$.
We will refer to $\Lambda_{i}$ as a Lagrangian cell. Using the partial ordering of the coisotropic cells (and any refinement of it) we obtain closed Lagrangian subvarieties $\Lambda_{\geq i}=\bigsqcup_{j\geq i}\Lambda_{j}$ and $\Lambda_{>i}=\bigsqcup_{j>i}\Lambda_{j}$. Note that under our assumptions that the closures of the $\Lambda_{i}$ are the irreducible components of $\Lambda$. The next lemma shows that a Lagrangian skeleton is compatible with the collection of symplectic reductions $\pi_{i}\colon C_{i}\to S_{i}$.
@@
If $(\mathfrak{X},\omega)$ is elliptic symplectic, and $\Lambda$ is a closed $\mathbb{G}_{m}$-invariant Lagrangian subvariety such that $\Lambda\cap C_{i}=\Lambda_{i}$ is smooth, then $\bar{\Lambda}_{i}=\pi_{i}(\Lambda_{i})$ is a smooth Lagrangian in $S_{i}$ and $\Lambda_{i}=\pi_{i}^{-1}(\bar{\Lambda}_{i})$.
@@
We must check that $\Lambda_{i}=\pi_{i}^{-1}(\pi_{i}(\Lambda_{i}))$, and that $\bar{\Lambda}_{i}=\pi_{i}(\Lambda_{i})$ is Lagrangian in $S_{i}$. This can be checked locally on $S_{i}$, and indeed since it is shown in [BDMN, §2] that $\pi_{i}$ factors the projection map $p_{i}\colon C_{i}\to Y_{i}$, it is enough to check this locally on $Y_{i}$. We thus fix $y\in Y_{i}$. Now $p_{i}$ is an affine bundle over $Y_{i}$, and by [BB] we may equivariantly trivialize it near $y$, so that we may assume $C_{i}=Y_{i}\times V$ where $V$ is a linear $\mathbb{G}_{m}$-representation with strictly negative weight. If $U<V$ denotes the subspace of vectors with weight $k$ where $-\ell\leq k<0$ and $W$ denotes the subspace of $V$ consisting of vectors with weights strictly less than $-\ell$, then it is shown in Lemma 2.5 of [BDMN] that in this trivialization we have $S=Y\times U$ where $\pi_{i}$ is the projection along $W$.
@@ 2.2. Better…Stronger…Faster
Now, suppose that, in addition, $\mathfrak{X}$ comes equipped with a $\mathsf{T}={\mathbb{G}}_{m}^{2}$ action, such that:
@@
2.3. The Lagrangian $\mathsf{\Lambda}$
Suppose next that $\mathfrak{X}$ is bionic symplectic. Let $\mathsf{H}=\mathsf{H}(s)\colon{\mathbb{G}}_{m}\rightarrow\operatorname{Aut}(\mathfrak{X})$ denote the Hamiltonian ${\mathbb{G}}_{m}$-action. We define the Lagrangian skeleton $\mathsf{\Lambda}$ of $\mathfrak{X}$ to be:
(2.1) $\mathsf{\Lambda}=\Big{\{}x\in\mathfrak{X}\,\Big{|}\,\lim_{s\rightarrow\infty}\mathsf{H}(s)\cdot x\,\,\text{exists in $\mathfrak{X}$}\Big{\}}.$
@@ 2.3. The Lagrangian $\mathsf{\Lambda}$
The subset $\mathsf{\Lambda}$ is a Lagrangian subset of $\mathfrak{X}$.
For each $i$, let $\mathsf{\Lambda}_{i}=C_{i}\cap\mathsf{\Lambda}$, where $C_{i}$ is a coisotropic cell as in Section 2.1.
@@
    Each $\mathsf{\Lambda}_{i}$ is a Lagrangian vector subspace of the ${\mathbb{G}}_{m}$-representation $C_{i}$.
@@
    Moreover, the projection $\pi_{i}(\mathsf{\Lambda}_{i})\subset S_{i}$ is a $\mathsf{T}$-stable Lagrangian subspace of the symplectic vector space $S_{i}$, and $\mathsf{\Lambda}_{i}=\pi^{-1}\big{(}\pi(\mathsf{\Lambda}_{i})\big{)}$.
@@
The variety $\mathsf{\Lambda}$ is a Lagrangian skeleton of $\mathfrak{X}$.
@@ 2.4. Cotangent Bundles of Spherical Varieties
We may write $X$ as the disjoint union $X=\bigsqcup_{s\in\mathcal{S}}\mathcal{O}_{s}$ of its $B$-orbits, where $|\mathcal{S}|<\infty$. To show that $T^{*}X$ has the structure of a bionic symplectic variety, it suffices to exhibit a $1$-parameter subgroup $\gamma\colon\mathbb{G}_{m}\to B$ such that $\gamma(\mathbb{G}_{m})$ has finitely many fixed points on $\mathcal{O}_{s}$ for each $s\in\mathcal{S}$. Indeed assuming such a $\gamma$ exists, it has an induced action on $T^{*}X$, and by modifying this action by a sufficiently high power of the scaling action on the fibres of $T^{*}X$ we may ensure that the resulting action on $T^{*}X$ is fibrewise convergent.
@@
If $X$ is either (a) toric, or (b) a partial flag manifold $G/P$, then $T^{*}X$ has a $\mathsf{T}$-action as above.
@@ 2.5. Hilbert Schemes of Points on Minimal Resolutions of Type $A$ Kleinian Singularities
Another example of a bionic symplectic variety is the $\mathsf{T}$-action on $({\mathbb{C}}^{2})^{[n]}$ induced from the action on ${\mathbb{C}}^{2}$. More generally, the action on the Hilbert scheme of points on the minimal resolution of the type $A$ Kleinian singularity is bionic. It is explained in [Go] how to identify the Hamiltonian $\mathbb{G}_{m}$ action and show that it has finitely many fixed points.
The algebra of invariant (with respect to the elliptic $\mathbb{G}_{m}$-action) global sections of a DQ-algebra on this Hilbert scheme is isomorphic to the spherical subalgebra of a rational Cherednik algebra associated to the wreath product group $\mathfrak{S}_{n}\wr\mu_{\ell}$. When localization holds, geometric category $\mathcal{O}$ is equivalent to the usual (spherical) category $\mathcal{O}$ introduced in [GGOR].

### v2

@@ 1. Introduction
Many symplectic varieties that arise in nature can be equipped not only with an elliptic ${\mathbb{G}}_{m}$-action but also a compatible Hamiltonian ${\mathbb{G}}_{m}$-action, i.e. one preserving the symplectic form. When the two actions have the same, finite, set of fixed points, we call the variety $\mathfrak{X}$ bionic symplectic: a precise definition appears in Section 2.2. Standard examples are, again, cotangent bundles of projective spherical varieties, and Hilbert schemes of points on minimal resolutions of cyclic Kleinian singularities (among other quiver varieties). Such a variety comes equipped with a canonical Lagrangian “skeleton” $\Lambda^{c}$.
Suppose now that a symplectic variety $\mathfrak{X}$ comes equipped with a deformation quantization ${\mathcal{W}}$ compatible with an elliptic ${\mathbb{G}}_{m}$-action. Taking ${\mathcal{W}}$-modules equivariant for the elliptic ${\mathbb{G}}_{m}$-action, one arrives at a category useful for geometric representation theory. When the action is the scaling action on fibers of $\mathfrak{X}=T^{*}X$, this category reproduces a category of twisted ${\mathcal{D}}$-modules on $X$; more generally such categories are closely related (depending on the geometry of $\mathfrak{X}$) to representations of Cherednik algebras, finite $W$-algebras, and more. When the variety is bionic symplectic and the deformation quantization is compatible with both ${\mathbb{G}}_{m}$-actions, there is a good geometric analogue of category $\mathcal{O}$ [BLPW1, BLPW2, Lo1, Web] inside the category of ${\mathcal{W}}$-modules, consisting of modules supported on the skeleton $\Lambda^{c}$.
The present paper establishes two basic features of the category of ${\mathcal{W}}$-modules on a bionic symplectic variety. We build a sheaf of proper dg algebras $\Omega$ that plays the role for ${\mathcal{W}}$-modules analogous to that played by the de Rham complex for ${\mathcal{D}}_{X}$-modules, i.e. ${\mathcal{W}}$-modules on $\mathfrak{X}=T^{*}X$. The sheaf $\Omega$ is supported (cohomologically) on a Lagrangian $\Lambda\supset\Lambda^{c}$ in $\mathfrak{X}$; while for general bionic symplectic varieties we do not have an explicit description of $\Lambda$, in the case where $\mathfrak{X}$ is a symplectic resolution (which is the case in essentially all examples of import to geometric representation theory), we have $\Lambda=\Lambda^{c}$ (c.f. 1.4 below).
@@
Assume that locally free resolutions exist for coherent ${\mathcal{W}}$-modules. Then the bounded derived category of coherent ${\mathcal{W}}$-modules is equivalent to the derived category of coherent dg $\Omega$-modules:
@@
As in [BDMN], we define the category of ”quasi-coherent ${\mathcal{W}}$-modules” ${\mathcal{W}}\text{-}{\mathsf{qcoh}}$ to be the ind-category of good ${\mathcal{W}}$-modules. We show that Theorem 1.1 extends to the derived category $D({\mathcal{W}}\text{-}{\mathsf{qcoh}})$, though not in the obvious way. We introduce the ”exotic” derived category $D_{ex}(\Omega\text{-}{\mathsf{Mod}})$ as a certain quotient of $K(\Omega\text{-}{\mathsf{Mod}})$, admitting a quotient functor $D_{ex}(\Omega\text{-}{\mathsf{Mod}})\to D(\Omega\text{-}{\mathsf{Mod}})$ to the usual derived category of dg $\Omega$-modules.
@@ 1.1. Symplectic resolutions
Assume that $\mathfrak{X}\to X$ is a symplectic resolution. Then the sheaf of dg algebras $\Omega$ is cohomologically supported on the canonical Lagrangian skeleton $\Lambda^{c}$.
The key to Theorem 1.4 is the fact that one can choose a local generator $\mathcal{L}$ (whose dg endomorphism ring is $\Omega$) supported on $\Lambda^{c}$; see Proposition 3.16.
1.2. Extension of holonomic complexes
@@
${D^{b}_{\mathrm{Hol},C}({\mathcal{W}}_{\mathfrak{X}})}$${D^{b}_{\mathrm{Hol}}({\mathcal{W}}_{\mathfrak{X}})}$${D^{b}_{\mathrm{Hol}}({\mathcal{W}}_{U}).}$$\scriptstyle{i_{*}}$$\scriptstyle{\mathbb{R}i^{!}}$$\scriptstyle{i^{*}}$$\scriptstyle{j^{*}}$$\scriptstyle{\mathbb{R}j_{*}}$$\scriptstyle{j_{!}}$
If $S$ denotes the coisotropic reduction of $C$ then it is shown in [BDMN, Theorem 1.7(2)] that the functor of Hamiltonian reduction $\mathbb{H}$ identifies $D^{b}_{\mathrm{Hol},C}({\mathcal{W}}_{\mathfrak{X}})$ with $D^{b}_{\mathrm{Hol}}({\mathcal{W}}_{S})$.
@@ 1.3. Categorical conventions
Acknowledgments
The authors are grateful to D. Ben-Zvi, K. Kremnitzer, J. Lebowski, D. Nadler, and D. Treumann for inspiration and helpful conversations. We would also like to thank the two referees who read the first version of the paper and provided many helpful comments, which undoubtedly improved the exposition in many places.
@@ 2.1. Lagrangian Skeletons
    For each $i\in I$, $\Lambda_{i}=\Lambda\cap C_{i}$ is a nonempty smooth closed subvariety of $C_{i}$.
We will refer to $\Lambda_{i}$ as a Lagrangian cell. Using the partial ordering of the coisotropic cells (and any refinement of it) we obtain closed Lagrangian subvarieties $\Lambda_{\geq i}=\bigsqcup_{j\geq i}\Lambda_{j}$ and $\Lambda_{>i}=\bigsqcup_{j>i}\Lambda_{j}$. Note that under our assumptions the closures of the $\Lambda_{i}$ are the irreducible components of $\Lambda$. The next lemma shows that a Lagrangian skeleton is compatible with the collection of symplectic reductions $\pi_{i}\colon C_{i}\to S_{i}$.
@@
If $(\mathfrak{X},\omega)$ is elliptic symplectic, and $\Lambda$ is a closed $\mathbb{G}_{m}$-invariant Lagrangian subvariety such that $\Lambda\cap C_{i}=\Lambda_{i}$ is smooth, then $\pi_{i}(\Lambda_{i})$ is a smooth Lagrangian in $S_{i}$ and $\Lambda_{i}=\pi_{i}^{-1}(\pi_{i}({\Lambda}_{i}))$.
@@
We must check that $\Lambda_{i}=\pi_{i}^{-1}(\pi_{i}(\Lambda_{i}))$, and that $\pi_{i}(\Lambda_{i})$ is Lagrangian in $S_{i}$. This can be checked locally on $S_{i}$, and indeed since it is shown in [BDMN, §2] that $\pi_{i}$ factors the projection map $p_{i}\colon C_{i}\to Y_{i}$, it is enough to check this locally on $Y_{i}$. We thus fix $y\in Y_{i}$. Now $p_{i}$ is an affine bundle over $Y_{i}$, and by [BB] we may equivariantly trivialize it near $y$, so that we may assume $C_{i}=Y_{i}\times V$ where $V$ is a linear $\mathbb{G}_{m}$-representation with strictly negative weight. If $U<V$ denotes the subspace of vectors with weight $k$ where $-\ell\leq k<0$ and $W$ denotes the subspace of $V$ consisting of vectors with weights strictly less than $-\ell$, then it is shown in Lemma 2.5 of [BDMN] that in this trivialization we have $S_{i}=Y_{i}\times U$, where $\pi_{i}$ is the projection along $W$.
@@ 2.2. Better…Stronger…Faster
Now, suppose that $\mathfrak{X}$ comes equipped with a $\mathsf{T}={\mathbb{G}}_{m}^{2}$ action, such that:
@@
Remark 2.6.
When one has a bionic symplectic variety, there is a bit of freedom in the choice of the torus $\mathsf{T}$. More precisely, this definition really only requires the action of the product $\mathbb{G}_{m}\times\mathbb{G}_{m}$ of the tori giving the elliptic and Hamiltonian actions. This torus will be isogenous to $\mathsf{T}$, but may not be equal to it; for our purposes replacing $\mathsf{T}$ by this torus would be completely harmless. However, the natural torus actions occurring in the examples below often use a $\mathsf{T}$ which is not simply the product of the tori giving the elliptic and Hamiltonian actions, so it seemed sensible to allow this extra flexibility in the definition.
@@
2.3. The Lagrangian $\Lambda^{c}$
Suppose that $\mathfrak{X}$ is bionic symplectic. Let $\mathsf{H}=\mathsf{H}(s)\colon{\mathbb{G}}_{m}\rightarrow\operatorname{Aut}(\mathfrak{X})$ denote the Hamiltonian ${\mathbb{G}}_{m}$-action. We define the canonical Lagrangian skeleton $\Lambda^{c}$ of $\mathfrak{X}$ to be:
(2.1) $\Lambda^{c}=\Big{\{}x\in\mathfrak{X}\,\Big{|}\,\lim_{s\rightarrow\infty}\mathsf{H}(s)\cdot x\,\,\text{exists in $\mathfrak{X}$}\Big{\}}.$
@@ 2.3. The Lagrangian $\Lambda^{c}$
The subset $\Lambda^{c}$ is a Lagrangian subset of $\mathfrak{X}$.
For each $i$, let $\Lambda^{c}_{i}=C_{i}\cap\Lambda^{c}$, where $C_{i}$ is a coisotropic cell as in Section 2.1.
@@
    Under the isomorphism $C_{i}\cong\mathbb{A}^{d_{i}}$ given above, each $\Lambda^{c}_{i}$ is a vector subspace of the ${\mathbb{G}}_{m}$-representation $C_{i}$.
@@
    Moreover, the projection $\pi_{i}(\Lambda^{c}_{i})\subset S_{i}$ is a $\mathsf{T}$-stable Lagrangian subspace of the symplectic vector space $S_{i}$, and $\Lambda^{c}_{i}=\pi^{-1}\big{(}\pi(\Lambda^{c}_{i})\big{)}$.
@@
The variety $\Lambda^{c}$ is a Lagrangian skeleton of $\mathfrak{X}$.
Example 2.10.
An illustrative example is provided by the minimal resolution of the Kleinian singularity of type $A$; this example is discussed further in 2.5 and 3.17 below.
Let $Y_{n}$ denote the Kleinian singularity of type $A_{n}$ and let $X_{n}$ be its minimal resolution, a smooth symplectic surface. (The symplectic form is induced from volume form on ${\mathbb{C}}^{2}$). There is a natural two-torus $\mathsf{T}$ acting on $X_{n}$, and it has $n$ fixed points $\{p_{1},p_{2},\ldots,p_{n}\}$ all lying in the exceptional fibre of the resolution: indeed if the exceptional locus is $D_{1}\cup D_{2}\ldots\cup D_{n-1}$ where each $D_{i}$ is a projective line then $D_{i}\cap D_{i+1}=\{p_{i+1}\}$ for $21\leq i\leq n-$ 2 and the other fixed points $p_{1}$ and $p_{n}$ lie on $D_{1}\backslash\{p_{1}\}$ and $D_{n-1}\backslash\{p_{n-1}\}$ respectively.
The Hamiltonian $\mathbb{G}_{m}$-action $\mathsf{H}\colon\mathbb{G}_{m}\to\mathsf{T}$ is induced from a Hamiltonian action on ${\mathbb{C}}^{2}$, and the locus of points $\Lambda^{c}$ which have a limit under it is the union of the exceptional locus and an affine line tangent to $C_{n-1}$ at $p_{n}$. On the open cell, one of course already has a symplectic variety, so it is its own reduction, while on each of the other cells, the reduction is just a point.
@@ 2.4. Cotangent Bundles of Spherical Varieties
We may write $X$ as the disjoint union $X=\bigsqcup_{s\in\mathcal{S}}\mathcal{O}_{s}$ of its $B$-orbits, where $|\mathcal{S}|<\infty$. To show that $T^{*}X$ has the structure of a bionic symplectic variety, it suffices to exhibit a $1$-parameter subgroup $\gamma\colon\mathbb{G}_{m}\to B$ such that $\gamma(\mathbb{G}_{m})$ has finitely many fixed points on $\mathcal{O}_{s}$ for each $s\in\mathcal{S}$. Indeed assuming such a $\gamma$ exists, it induces a Hamiltonian action on $T^{*}X$, and by modifying this action by a sufficiently high power of the scaling action on the fibres of $T^{*}X$ we may ensure that the resulting action on $T^{*}X$ is fibrewise convergent.
@@
If $X$ is either a partial flag manifold or a toric variety then $T^{*}X$ is bionic symplectic.
@@ 2.5. Hilbert Schemes of Points on Minimal Resolutions of Type $A$ Kleinian Singularities
Another example of a bionic symplectic variety is the $\mathsf{T}$-action on $({\mathbb{C}}^{2})^{[n]}$ induced from the action on ${\mathbb{C}}^{2}$. More generally, the action on the Hilbert scheme of points on the minimal resolution of the type $A$ Kleinian singularity is bionic. It is explained in [Go] how to identify the Hamiltonian $\mathbb{G}_{m}$-action and show that it has finitely many fixed points.
The algebra of invariant (with respect to the elliptic $\mathbb{G}_{m}$-action) global sections of a DQ-algebra on this Hilbert scheme is isomorphic to the spherical subalgebra of a rational Cherednik algebra associated to the wreath product group $\mathfrak{S}_{n}\wr\mu_{\ell}$; see [Go06] for a proof of this. When localization holds, geometric category $\mathcal{O}$ is equivalent to the usual (spherical) category $\mathcal{O}$ introduced in [GGOR].